OpenStudy (anonymous):

Please help me solve the following sum on complex numbers:

OpenStudy (anonymous):

GIven that (3+2i) is a root of \[2x^3 + px^2 + 20x + q = 0\] and p and q are real numbers, find p and q (finding the other roots is the next part of the question which I could do much more easily, no worries :P)

OpenStudy (experimentx):

put x = 3 + 2i, and simplify simplify ... and compare the real and imaginary part of both sides.

OpenStudy (anonymous):

oh whoops, thank you so much :D

OpenStudy (experimentx):

and for other roots, factor out (x - 3 - 2i) <-- this can be quite troublesome.

OpenStudy (anonymous):

see, one of the roots has to be 3-2i, right?

OpenStudy (experimentx):

oh ,, yes yes ... right, it's cubic equation.

OpenStudy (anonymous):

then we multiply those two, and divide?? idk, lemme work it out and see.

OpenStudy (experimentx):

best way is to guess the root by remainder theorem. else you can factor out (x^2 - 6x + 13)

OpenStudy (anonymous):

what do you mean by compare the imaginary and real things on both sides??

terenzreignz (terenzreignz):

Just set x = 3+2i for now... and simplify.

OpenStudy (anonymous):

I did that. The previous part of the question said find x^2 and x^3 if x = 3+2i. I'm wondering, maybe i've got that wrong

terenzreignz (terenzreignz):

\[\Large 2(3+2i)^3+\color{red}p(3+2i)^2+20(3+2i)+\color{blue}q=0\]

OpenStudy (anonymous):

OH MY GOD. I JUST REALIZED

terenzreignz (terenzreignz):

What IS the cube and square of 3+2i?

OpenStudy (anonymous):

so instead of multiplying normally, I went and used de moivres theorem, and got some really weird answers

terenzreignz (terenzreignz):

de moivre? You actually used polar forms? -.-

OpenStudy (anonymous):

waste of time, and energy, and brain power. UGH

terenzreignz (terenzreignz):

So... I take it you get it now? ^_^

terenzreignz (terenzreignz):

Hey... everything okay?